# Forgetting-MarI: LLM Unlearning via Marginal Information Regularization

### v2

@@ E.1 Detector Methods
  • White-box methods: Perplexity outlier and reference model perplexity outlier [carlini2021extract], domain normalized minimum k-percentage [zhang2024dcpdd], and data-set level inference [maini2024dataset].
@@
  • Black-box methods: Direct regurgitation probes [karamolegkou2023leak], Name-cloze membership inference [chang2023cloze], DE-COP: multi-choice preference [duarte2024decop], Output-consistency measures [dong2024consistency], and VeilProbe [hu2024veilprobe].
@@
Black-box methods are weaker detectors than white-box methods since they do not have access to a model’s internals. Since our method assumes access to the model parameter, we tested our method against the current SotA white-box method, the minimum k-percent method [shi2024mink], to demonstrate the effectiveness of our unlearning in real-world applications.

### v3

@@ E.1 Detector Methods
  • White-box methods: Perplexity outlier and reference model perplexity outlier [3], domain normalized minimum k-percentage [40], and data-set level inference [30].
@@
  • Black-box methods: Direct regurgitation probes [21], Name-cloze membership inference [9], DE-COP: multi-choice preference [5], and Output-consistency measures [16].
@@
Black-box methods are weaker detectors than white-box methods since they do not have access to a model’s internals. Since our method assumes access to the model parameter, we tested our method against the current SotA white-box method, the minimum k-percent method [34], to demonstrate the effectiveness of our unlearning in real-world applications.